OpenStudy (anonymous):

is there anyone in algebra 1 foundations b ?

OpenStudy (mathlegend):

What is the question @slanginchoppas ?

OpenStudy (anonymous):

I can help you if you post questions, i'm not going to do you're entire assignment for you. but i can help.

OpenStudy (anonymous):

what is the length of the hypotenuse of the right triangle shown ? round your answers if needed . one side is 6 , one side is 11

OpenStudy (anonymous):

thats simple: here is your equation\[a^2 + b^2 = c^2\]

OpenStudy (anonymous):

You have a and b because C is always the hypotenuse

OpenStudy (anonymous):

what you're going to do is get 6(Squared) and 11(squared) and add them together after you add them together you find the square root of the sum so 36 + 121 = 157 and then you find the square root of 157

OpenStudy (anonymous):

thankyou !!! :)

OpenStudy (anonymous):

of course friend.

OpenStudy (anonymous):

what about if you need to find B ?

OpenStudy (anonymous):

Like I have the hypotenuse & A , the side that I need to find is B

OpenStudy (anonymous):

Then you would have to move by itself so you would subtract a and put it with the hypotenuse.

OpenStudy (anonymous):

a2 + b2 = c2 you would subtract a2 so it would be c2-a2=b2